# Nordlund0000-0001-6244-1942

### v1

@@ II Quantum chemistry of repulsive interatomic potentials
The interaction of bare nuclei without any electrons (in the absence of nuclear resonances, which are beyond the scope of this work), is given by the Coulomb potential, expressed in standard units as
$V_{\rm Coul}(r)={e^{2}\over 4\pi\varepsilon_{0}}{Z_{1}Z_{2}\over r}$ (1)
where the constant $e^{2}/(4\pi\varepsilon_{0})$ equals $14.000\,\mathrm{eV\AA}$, and $Z_{1}$ and $Z_{2}$ are the atomic numbers of the colliding nuclei.
@@
The two features discussed above make the screening function attractive for use in simulations: unlike the bare interatomic potential V(r), $\phi(r)$ has a limited numerical range and varies smoothly across all internuclear distances, making it easy to interpolate it accurately from a moderate number of tabulated points, for example. It is worth noting that also chemical interactions can be considered in this formalism, since the mapping $r\to V_{\rm Coul}(r)$ is invertible, and it is easy to see that the description of a bonded diatomic molecule leads to $\phi(r)<0$ at chemically bound interaction distances $r$ [12]. Antiproton-atom interactions can also be modeled in this formalism [21], even though the Coulomb interaction between the nuclei is then attractive instead of repulsive.
@@
To understand this, let us consider a simple thought experiment: when $r\to 0$, the electron density of the system of two atoms $Z_{1}$ and $Z_{2}$ must in fact approach the electron density of the united atom with atomic number $Z_{1}+Z_{2}$. This case is achieved in practice already when the internuclear distance is smaller than the $1s$ orbital of the united atom; we denote this distance $r_{\rm ua}$, where “$\rm ua$” stands for united atom. If superposition could be applied when $Z_{1}=Z_{2}=Z$, the electron density of the united atom with atomic number $2Z$ would be simply twice that of $Z$. Yet, already the analysis of the hydrogenic atom shows that this is not true: the radius of the $1s$ orbital of the united atom is actually half that of the separate atoms. As screening effects are negligible for the $1s$ orbital, this simple argument shows that the superposition of atomic densities leads to a completely incorrect form for the density at the united atom limit.
@@
With a correct description of the united atom limit, the interatomic potential can be reliably computed at small $r$, $r\lesssim r_{\rm ua}$, using perturbation theory [37], for example. However, capturing the electronic structure at intermediate separations, $r_{\rm ua}<r\ll r_{\rm equi}$, where the inner electronic shells of the two atoms partly overlap while the outermost shells resemble those of the united atom [38], is quite challenging. Excluding the case of hydrides, where the united atom model yields a qualitatively correct model of the electronic structure of the diatomic molecule [39, 40], in the general case the orbitals of the diatomic molecule undergo significant relaxation effects from those of the free atoms, and the chemistry of the high-energy region remains relatively unexplored.
@@
As already mentioned above, we here examine three ways to calculate repulsive interatomic potentials from quantum chemical methods, throughout the three distance regimes discussed above. Previous work in Ref. 43 has shown that the error in $\phi(r)$ is almost totally governed by the quality of the one-particle basis set, with the differences between various levels of theory (e.g. Hartree–Fock vs coupled-cluster theory vs density functional theory) being negligible with respect to the pursued level of accuracy. Similarly, the role of the employed charge state is expected to have a negligible effect on the repulsive barrier, as differences of the order of 1 eV, that would be unacceptable for modeling the chemical equilibrium, become negligible in the highly repulsive part of the interatomic potential dominated by $V_{\rm Coul}(r)$.
As most electronic structure methods have been developed for the region $r\approx r_{\rm equi}$, a major problem in the present effort is to find a stable numerical representation that can adapt to the varying electronic configurations found in the regime $r_{\rm ua}<r<r_{\rm equi}$. A breakthrough was recently achieved in Ref. 41 in showing that the strongly repulsive region can be faithfully modeled with established atomic basis set approaches, if the united atom limit is included in the construction of the basis set and if the significant linear dependencies that arise in the basis set at small internuclear distances are properly taken care of. We will employ this method to verify the accuracy of the calculations carried out with the two other methods discussed below.
@@ III Methods
III.1 Gaussian-type orbital calculations
Preceding work in Ref. 43 found that the repulsive potential can be accurately reproduced with Gaussian-type orbital (GTO) calculations, if additional basis functions are added on the bond between the nuclei. In this work, we follow Ref. 41 and go beyond the use of bond functions by ensuring that the united atom limit is correctly reproduced by the calculation. With this aim, we introduce ghost basis functions [46] corresponding to the united atom at the center of charge at $z(r)=Z_{B}r/(Z_{A}+Z_{B})$, when atoms $A$ and $B$ are placed at the origin and at $(0,0,r)$, respectively. State-of-the-art polarization consistent (pc-$n$) GTO basis sets by [47] are employed in all calculations, with the basis set for the ghost atom also including additional polarization functions. To allow even the deep core orbitals to hybridize, i.e., to form bonding and antibonding orbitals, all the GTO basis sets are employed in fully uncontracted form. After testing, the final results are shown for the triple-$\zeta$ pc-2 basis set, which is expected to afford screening functions of high quality.
The resulting molecular basis set has pathological linear dependencies for vanishingly small interatomic distance, $r\to 0$. A pivoted Cholesky decomposition procedure was employed to remove linearly dependent basis functions [48]; this method was shown in Ref. 41 to afford excellent accuracy in Hartree–Fock (HF) calculations compared to ones performed with the HelFEM program [49, 50] that employs numerical basis functions following the finite element method, reproducing values directly at the complete basis set limit.
@@ III.1 Gaussian-type orbital calculations
III.2 Numerical atomic orbital calculations
@@
III.3 Orbital-free density functional calculations
@@ IV.1 Comparison with earlier model calculations
The issue of the basis set truncation error was revisited in the recent study of Lehtola [41], who showed that an enriched basis similar to the one employed in this work (see section III.1) affords an excellent level of agreement with fully numerical calculations at the Hartree–Fock level of theory. In the following, we therefore assume that the MP2 calculations carried out in the fully uncontracted triple-$\zeta$ pc-2 basis set enriched with united-atom basis functions at the center of nuclear charge (see section III.1) provide a good estimate of the screening functions.
We begin our analysis by comparing the new MP2 results with the two versions of the ZBL potential, and with earlier results of Nordlund et al. [43], both their fully numerical Hartree–Fock calculation and their original DMol calculations. Results for the Si-Si system are illustrated in fig. 2, where the top part shows the potential energy curve $V(r)$, the middle part shows the screening function $\phi(r)$, and the bottom part shows the ratio of the screening functions. Since the differences between the methods are nearly undiscernable in the top and middle plots, we also plot the ratios of the resulting potential energies in the bottom part of the plot to make the differences better visible.
The results in fig. 2 show that the ZBL universal potential starts deviating from the pair-specific quantum chemical approaches already above 0.02 Å, whereas the quantum chemical approaches agree well with each other up to a distance of 0.3 Å. In terms of potential energy, these distances correspond to about 100 keV and 2 keV, respectively. It is remarkable that even though the purely repulsive ZBL potential is specifically meant for the high-energy regions, it still starts deviating from the quantum chemical data already at energies below 100 keV.
@@ IV.2 Comparison of selected potentials
Figure 3 compares screening functions for four atom pairs: H-H, H-Si, Si-Si and Ti-Si. In the case of H-H, the ZBL potential clearly deviates from the quantum chemical calculations (DMol and MP2). The quantum chemical methods agree well with each other within the plotting accuracy, in line with earlier findings of Refs. 41 and 43.
@@ IV.3 Comparison to orbital-free calculations
is shown in fig. 8 for the studied set of homoatomic systems ($Z_{1}=Z_{2}=Z$ in eq. 10). As the near-equilibrium attractive part of the potential is not of interest in this work, the average in eq. 10 was taken over the subset of points for which the potential is either larger than 10 eV or 100 eV; this also ensures that $\phi(r_{i})>0$ in eq. 10. The resulting mean absolute errors for the two cutoffs are 2.7% and 1.3%, respectively, confirming that the OF-DFT model yields similar screening functions regardless of whether the atomic electron density is taken from gas-phase or solid-state calculations.
@@ V Analytical potential fit
$V_{\rm NLH}(r)={1\over 4\pi\varepsilon_{0}}{Z_{1}Z_{2}\over r}\phi_{\rm NLH}(r),$ (13)
which enables facile evaluation of forces and Hessians, for instance.
@@
Figure 11: Average difference $\Delta_{Z_{1},Z_{2}}$ between the NLH and the MP2 potentials for all cases within the nonrelativistic limit. The analysis was done at distances where the MP2 potential is either above 10, 30 or 100 eV.
Figure 12: Average $\bar{\Delta}$ and root-mean-square deviation $\sigma_{\Delta}$ of the NLH potential fit over all element pairs. All numbers are given in %. Note that the ordinate scale is 10 times smaller than in fig. 10.
@@ VI.1 Comparison of potentials with each other
The results of these calculations are shown in fig. 13 for Si-Si implantation, in fig. 14 for Ar-Al implantation, in fig. 15 for H-Si implantation, and in fig. 16 for Fe-Fe implantation. A summary of the results is given in table 4. The choice of potential clearly matters: there are differences in the range profiles as well as mean ranges between the ZBL and quantum chemical DMol and MP2 potentials that are comparable to differences in the interatomic potentials (of the order of $5\%$).
@@ H Si 10 Non-channeled 1753 $\pm$ 2 1751 $\pm$ 2 1758 $\pm$ 2 1750 $\pm$ 2 1750 $\pm$ 2
To get more insight on whether atomic size affects the sensitivity of the range calculation on the interatomic potential, we simulated ions in the whole studied range $1\leq Z\leq 92$ implanted in the same [110] channeling direction of Si for which a clear sensitivity was observed in the case of H. This comparison was done for mean ranges obtained with the DMol and NLH potentials. Since the latter potential is a fit to the DMol data above 30 eV, they agree within $\sim$ 2% at high energies (see section IV.4). However, the NLH and DMol potentials are quite different at low energy since the NLH potential is repulsive at all internuclear distances $r$, while the DMol potential has an attractive part at distances close to the chemical equilibrium. Data in the inset of fig. 15 shows that there are sometimes large differences in the mean ranges for small atomic numbers $Z\lesssim$ 10. For larger atomic numbers, the difference is at most $\sim$ 2%, which may be explained by the higher interaction energies of heavier ions. Hence, the results indicate that small atoms of low-atomic number may be sensitive to the choice of attractive potential, when implanted in wide channels such as the [110] channel in Si. Comparisons of different low-energy interaction parts using the data provided in the Supplementary Information can be used to test whether a particular channeling condition is indeed sensitive to low-energy interactions.
However, we note that using a pair potential with the attractive part of a diatomic system may also be misleading because at the low, near-equilibrium energies below 100 eV, the interatomic interactions are many-body in character. In other words, the attractive part of the potential in a solid depends on how many neighbours an atom has in a particular configuration [91, 92], and hence will be different from that of the diatomic molecule. The question on how to join repulsive potentials with equilibrium attractive ones is discussed in appendix A.
@@ VI.2 Comparison of range profiles with experiments
The experimental depth profiles were scanned in and digitized from Refs. 76, 93 with the g3data code [94]. From comparison of repeated digitizations, this process is estimated to introduce errors $\lesssim$ 1% to the data. The experimental data may carry larger systematic errors associated with the SIMS method itself. However, since the original references do not give estimates of experimental errors, error bars are not reported here for the experimental data.
Simulations were carried out with the same MDRANGE approach described earlier, except that a surface oxide layer of 2.2 nm thickness was included in the calculations; the thickness of the surface layer corresponds to 4 Si unit cells. The electronic stopping power was described using the density-functional based Puska–Echenique–Nieminen–Ritchie model utilizing a 3D electron density of Si [83], combined with a Firsov local stopping. This approach has previously been shown to give excellent agreement between experimental and simulated range profiles in Si [85].
For the case of As implantation into amorphous Si, the nonlocal TRIM 1996 electronic stopping was used [78]. This electronic stopping was also used for the amorphous 2.2 nm SiO2 surface layers. The amorphous Si was modelled with a 4 nm cube of random atom positions with a minimum separation of 2.1 Å. This approach has previously been shown to well describe implantation into amorphous material [59]. The ion-oxygen interactions were modelled with the ZBL universal repulsive potential; since the surface oxide layer is very thin, using any other potential would not make any significant difference. We verified this by simulating the case of 180 keV As implantation into amorphous Si also with the NLH potential for the As-O interactions. As expected, the mean ranges agreed within the statistical uncertainty of $\pm$ 1 Å.
@@
The simulations for which mean ranges are given in table 5 were carried out for 100 000 ions; the statistics is reflected in the error of the mean. The dose used in the simulations was the same as the experimental fluence of 1013 ions/cm2, except for the case of 200 keV P ions, for which it was 2.5×2.5$\times$2.5 ×1013 ions/cm2. For the cases shown in figs. 18 to 20, the simulations were carried out for one million ions for clarity in the plots. The comparisons with experiments contain no adjustable parameters, and both the depth and concentration scales are comparable in absolute values.
@@
Figure 18: Range profiles for As ions implanted into Si compared with experimental data [76]. The data labeled ”MD” are simulated with the MD-RIA approach with the MDRANGE code, and ”SRIM” with the commonly used BCA code SRIM.

### v3

@@ II Quantum chemistry of repulsive interatomic potentials
II.1 Screening function formalism for repulsive potentials
To understand the scientific basis of the repulsive potential calculations, we first discuss the approaches used including recent developments that lay the basis of the results of the paper. The interaction of bare nuclei without any electrons (in the absence of nuclear resonances, which are beyond the scope of this work), is given by the Coulomb potential, expressed in red international standard (SI) units as
$V_{\rm Coul}(r)={1\over 4\pi\varepsilon_{0}}{Z_{1}eZ_{2}e\over r}.$ (1)
Here $\varepsilon_{0}$ is the dielectric constant $=8.854187817\times 10^{-12}$ F/m, $Z_{1}$ and $Z_{2}$ are the atomic numbers of the two colliding atoms, $e$ the elemental charge $=1.602176634\times 10^{-19}$ C, and $r$ the distance between the two atoms.
@@ II.1 Screening function formalism for repulsive potentials
The two features discussed above make the screening function attractive for use in simulations: unlike the bare interatomic potential $V(r)$, $\phi(r)$ has a limited numerical range and varies smoothly across all internuclear distances, making it easy to interpolate it accurately from a moderate number of tabulated points, for example. It is worth noting that also chemical interactions can be considered in this formalism, since the mapping $r\to V_{\rm Coul}(r)$ is invertible, and it is easy to see that the description of a bonded diatomic molecule leads to $\phi(r)<0$ at chemically bound interaction distances $r$ [12]. Recently, one of us has shown that also antiproton-atom interactions can be modeled in this formalism [21], even though the Coulomb interaction between the nuclei is then attractive instead of repulsive.
@@ II.2 United atom approach for repulsive potentials
To understand this, let us consider a simple thought experiment: when $r\to 0$, the electron density of the system of two atoms $Z_{1}$ and $Z_{2}$ must in fact approach the electron density of the united atom with atomic number $Z_{1}+Z_{2}$. This case is achieved in practice already when the internuclear distance is smaller than the $1s$ orbital of the united atom; we denote this distance $r_{\rm ua}$, where “$\rm ua$” stands for united atom. This united atom approach has been recently introduced for repulsive potentials Ref. 37, and in the current work we apply and analyze it systematically for all atom pairs in the non-relativistic limit.
If superposition could be applied when $Z_{1}=Z_{2}=Z$, the electron density of the united atom with atomic number $2Z$ would be simply twice that of $Z$. Yet, already the analysis of the hydrogenic atom shows that this is not true: the radius of the $1s$ orbital of the united atom is actually half that of the separate atoms. As screening effects are negligible for the $1s$ orbital, this simple argument shows that the superposition of atomic densities leads to a completely incorrect form for the density at the united atom limit.
@@
With a correct description of the united atom limit, the interatomic potential can be reliably computed at small $r$, $r\lesssim r_{\rm ua}$, using perturbation theory [38], for example. However, capturing the electronic structure at intermediate separations, $r_{\rm ua}<r\ll r_{\rm equi}$, where the inner electronic shells of the two atoms partly overlap while the outermost shells resemble those of the united atom [39], is quite challenging. Excluding the case of hydrides, where the united atom model yields a qualitatively correct model of the electronic structure of the diatomic molecule [40, 41], in the general case, the orbitals of the diatomic molecule undergo significant relaxation effects from those of the free atoms, and the chemistry of the high-energy region remains relatively unexplored.
@@
Previous work in Ref. 43 has shown that the error in $\phi(r)$ is almost totally governed by the quality of the one-particle basis set, with the differences between various levels of theory (e.g. Hartree–Fock vs coupled-cluster theory vs density functional theory) being negligible with respect to the pursued level of accuracy. Similarly, the role of the employed charge state is expected to have a negligible effect on the repulsive barrier, as differences of the order of 1 eV, that would be unacceptable for modeling the chemical equilibrium, become negligible in the highly repulsive part of the interatomic potential dominated by $V_{\rm Coul}(r)$.
As most electronic structure methods have been developed for the region $r\approx r_{\rm equi}$, a major problem in the present effort is to find a stable numerical representation that can adapt to the varying electronic configurations found in the regime $r_{\rm ua}<r<r_{\rm equi}$. A major methodological advance was recently achieved by one of us Ref. 37, in showing that the strongly repulsive region can be faithfully modeled with established atomic basis set approaches, if the united atom limit is included in the construction of the basis set and if the significant linear dependencies that arise in the basis set at small internuclear distances are properly taken care of. In the current work, we employ this method to verify the accuracy of the calculations carried out with the two other methods discussed below.
@@ III Methods
III.1 Gaussian-type orbital calculations (”MP2”)
Preceding work in Ref. 43 found that the repulsive potential can be accurately reproduced with Gaussian-type orbital (GTO) calculations, if additional basis functions are added on the bond between the nuclei. In this work, we follow Ref. 37 and go beyond the use of bond functions by ensuring that the united atom limit is correctly reproduced by the calculation. With this aim, we introduce ghost basis functions [46] corresponding to the united atom at the center of charge at $z(r)=Z_{B}r/(Z_{A}+Z_{B})$, when atoms $A$ and $B$ are placed at the origin and at $(0,0,r)$, respectively. State-of-the-art polarization consistent (pc-$n$) GTO basis sets by [47] are employed in all calculations, with the basis set for the ghost atom also including additional polarization functions. To allow even the deep core orbitals to hybridize, i.e., to form bonding and antibonding orbitals, all the GTO basis sets are employed in fully uncontracted form. After testing, the final results are shown for the triple-$\zeta$ pc-2 basis set, which is expected to afford screening functions of high quality.
The resulting molecular basis set has pathological linear dependencies for vanishingly small interatomic distance, $r\to 0$. A pivoted Cholesky decomposition procedure was employed to remove linearly dependent basis functions [48]; this method was shown in Ref. 37 to afford excellent accuracy in Hartree–Fock (HF) calculations compared to ones performed with the HelFEM program [49, 50] that employs numerical basis functions following the finite element method, reproducing values directly at the complete basis set limit.
@@ III.1 Gaussian-type orbital calculations (”MP2”)
III.2 Numerical atomic orbital calculations (”DMol”)
@@
III.3 Orbital-free density functional calculations (”ZBL pair-specific”)
@@ IV.1 Comparison with earlier model calculations
The issue of the basis set truncation error was revisited in the recent study of Lehtola [37], who showed that an enriched basis similar to the one employed in this work (see section III.1) provides an excellent level of agreement with fully numerical calculations at the Hartree–Fock level of theory. In the following, we therefore assume that the MP2 calculations carried out in the fully uncontracted triple-$\zeta$ pc-2 basis set enriched with united-atom basis functions at the center of nuclear charge (see section III.1) provide a good estimate of the screening functions.
We begin our analysis by comparing the new MP2 results with the two versions of the ZBL potential, and with the earlier results of Nordlund et al. [43], both their fully numerical Hartree–Fock calculation and their original DMol calculations. Results for the Si-Si system are illustrated in fig. 2, where the top part shows the potential energy curve $V(r)$, the middle part shows the screening function $\phi(r)$, and the bottom part shows the ratio of the screening functions. Since the differences between the methods are nearly undiscernable in the top and middle plots, we also plot the ratios of the resulting potential energies in the bottom part of the plot to make the differences better visible.
The results in fig. 2 show that the universal potential of ZBL starts to deviate from the pair-specific quantum chemical approaches already above 0.02 Å, whereas the quantum chemical approaches agree well with each other up to a distance of 0.3 Å. In terms of potential energy, these distances correspond to about 100 keV and 2 keV, respectively. It is remarkable that even though the purely repulsive ZBL potential is specifically meant for the high-energy regions, it still starts deviating from the quantum chemical data already at quite small distances and high energies. E.g. at the distance of $r=0.1$ where the ZBL potential has a value of 13.7 keV, it already deviates by 4 % from the MP2 potential.
@@ IV.2 Comparison of selected potentials
Figure 3 compares screening functions for four atom pairs: H-H, H-Si, Si-Si and Ti-Si. In the case of H-H, the ZBL potential clearly deviates from the quantum chemical calculations (DMol and MP2). The quantum chemical methods agree well with each other within the plotting accuracy, in line with earlier findings of Refs. 37 and 43.
@@ IV.3 Comparison to orbital-free calculations
is shown in fig. 8 for the studied set of homoatomic systems ($Z_{1}=Z_{2}=Z$ in eq. 10). As the near-equilibrium attractive part of the potential is not of interest in this work, the average in eq. 10 was taken over the subset of points for which the potential is either larger than 10 eV or 100 eV; this also ensures that $\phi(r_{i})>0$ in eq. 10. The resulting mean absolute errors for the two cutoffs are 2.7% and 1.3%, respectively, confirming that the OF-DFT model yields similar screening functions regardless of whether the atomic electron density is taken from gas-phase or solid-state calculations. In Appendix A, a systematic approach to join diatomic pair potentials to equilibrium potentials based on solid state DFT data is presented. This joining procedure should, at least in principle, also correct for the lack of solid-state effects in the diatomic potentials.
@@ V Analytical potential (”NLH”) fit
$V_{\rm NLH}(r)={1\over 4\pi\varepsilon_{0}}{Z_{1}eZ_{2}e\over r}\phi_{\rm NLH}(r),$ (13)
which enables facile evaluation of forces, for instance. The screening function $\phi$ is unitless.
@@
Figure 11: Average difference $\Delta_{Z_{1},Z_{2}}$ between the NLH and the MP2 potentials for all cases within the nonrelativistic limit. The analysis was done at distances where the MP2 potential is either above 10, 30 or 100 eV. Note that the ordinate scale is 3 times smaller than in fig. 9
Figure 12: Average difference $\Delta_{Z_{1},Z_{2}}$ between the NLH and the DMol potentials for all cases within the nonrelativistic limit. The analysis was done at distances where the MP2 potential is either above 10, 30 or 100 eV. Note that the ordinate scale is 10 times smaller than in fig. 10.
@@ VI.1 Comparison of potentials with each other
The results of these calculations are shown in fig. 13 for the implantation of Si-Si, in fig. 14 for the implantation of Ar-Al, in fig. 15 for the implantation of H-Si, and in fig. 16 for the implantation of Fe-Fe. The depth distributions in the plots are area normalized to 1 to be in the form of a probability distribution of the implantation depth. A summary of the results is given in table 4. The error bars were calculated as the standard error of the mean, i.e. the standard deviation of the values of the individual ion range divided by the square root of the number of ions [91]. The choice of potential clearly matters: there are statistically significant differences in the range profiles as well as mean ranges between the ZBL and quantum chemical DMol and MP2 potentials that are comparable to differences in the interatomic potentials (of the order of $5\%$).
@@ H Si 10 Non-channeled 1753 $\pm$ 2 1751 $\pm$ 2 1758 $\pm$ 2 1750 $\pm$ 2 1750 $\pm$ 2
To get more insight on whether atomic size affects the sensitivity of the range calculation on the interatomic potential, we simulated ions in the whole studied range $1\leq Z\leq 92$ implanted in the same [110] channeling direction of Si for which a clear sensitivity was observed in the case of H. This comparison was done for mean ranges obtained with the DMol and NLH potentials. Since the latter potential is a fit to the DMol data above 30 eV, they agree within $\sim$ 2% at high energies (see section IV.4). However, the NLH and DMol potentials are quite different at low energy since the NLH potential is repulsive at all internuclear distances $r$, while the DMol potential has an attractive part at distances close to the chemical equilibrium. Data in the inset of fig. 15 shows that there are sometimes large differences in the mean ranges for small atomic numbers $Z\lesssim$ 10. For larger atomic numbers, the difference is at most $\sim$ 2%, which may be explained by the higher interaction energies of heavier ions. Hence, the results indicate that small atoms of low-atomic number may be sensitive to the choice of attractive potential, when implanted in wide channels such as the [110] channel in Si. Comparisons of different low-energy interaction parts using the data provided in the Supplementary Information [92] can be used to test whether a particular channeling condition is indeed sensitive to low-energy interactions.
However, we note that using a pair potential with the attractive part of a diatomic system may also be misleading because at the low, near-equilibrium energies below 100 eV, the interatomic interactions are many-body in character. In other words, the attractive part of the potential in a solid depends on how many neighbours an atom has in a particular configuration [93, 94], and hence will be different from that of the diatomic molecule. To account for such effects, one can use a many-body analytical or machine-learned potential that also accounts for chemical effects [95, 96, 94, 97, 98]. We present an approach by which one can join equilibrium potentials to the currently developed repulsive ones without any fitting parameters in Appendix A.
@@ VI.2.1 Simulation setup
The experimental depth profiles were scanned in and digitized from Refs. 76, 99 with the g3data code [100]. From comparison of repeated digitizations, this process is estimated to introduce errors $\lesssim$ 1% to the data. The experimental data may carry larger systematic errors associated with the SIMS method itself. However, since the original references do not give estimates of experimental errors, error bars are not reported here for the experimental data.
Simulations were carried out with the same MDRANGE approach described earlier, except that a surface oxide layer of 2.2 nm thickness was included in the calculations; the thickness of the surface layer corresponds to 4 Si unit cells. The sample atoms were given thermal displacements corresponding to 300 K using eq. 15, which gave a 1D root-mean-square displacement magnitude of 0.079 Å for the Si atoms. The electronic stopping power was described using the density-functional based Puska–Echenique–Nieminen–Ritchie model utilizing a 3D electron density of Si [83], combined with a Firsov local stopping. This approach has previously been shown to give excellent agreement between experimental and simulated range profiles in Si [85].
For the case of As implantation into amorphous Si, the nonlocal TRIM 1996 electronic stopping was used [78]. This electronic stopping was also used for the amorphous 2.2 nm SiO2 surface layers. The amorphous Si was modelled with a 4 nm cube of random atom positions with a minimum separation of 2.1 Å. This approach has previously been shown to well describe implantation into amorphous material [59]. The ion-oxygen interactions were modelled with the ZBL universal repulsive potential; since the surface oxide layer is very thin, using any other potential would not make any significant difference. We verified this by simulating the cases of 180 keV As implantation into amorphous Si and 15 keV As into Si in non-channeling direction also with the NLH potential for the As-O interactions. As expected, in both cases the mean ranges agreed within the statistical uncertainty of $\pm$ 1 Å.
@@
The simulations for which mean ranges are given in table 6 were carried out for 100 000 ions; the statistics is reflected in the standard error of the mean. At the bottom of the table, the average deviation for each potential is the arithmetic average over each individual deviation ($R_{\rm mean,sim,i}/R_{\rm mean,expt,i}-1$), where sim stands for the simulated and expt for the experimental values. The root-mean-square (rms) deviations were calculated as
$\textrm{rms}=\sqrt{\frac{\sum_{i=1}^{N}\left(R_{\rm mean,sim,i}/R_{\rm mean,expt,i}-1\right)^{2}}{N_{\rm pot}}}$ (17)
where $i$ loops over the various cases simulated (ion, energy, direction), and $N_{\rm pot}$ is the total number of cases simulated with this potential (10 for MP2 and 16 for the other potentials).
The dose used in the simulations was the same as the experimental fluence of 1013 ions/cm2, except for the case of 200 keV P ions, for which it was 2.5×2.5$\times$2.5 ×1013 ions/cm2. For the cases shown in figs. 18 to 20, the simulations were carried out for one million ions for clarity in the plots. The comparisons with experiments contain no adjustable parameters, and both the depth and concentration scales are comparable in absolute values.
@@
Figure 18: Range profiles for As ions implanted into Si compared with experimental data [76]. The data labeled ”MD” are simulated with the MD-RIA approach with the MDRANGE code, and ”SRIM” with the commonly used BCA code SRIM. The insets show how the range profile depends on the thickness of the amorphous oxide layer.
@@
VI.2.2 Effect of oxide thickness
To examine how sensitive the results are to the oxide layer thickness, we ran the case of 15 keV As implantation into the 100 channel and 100 keV As implantation into the non-channeling thickness with only 1 nm oxide thickness as well as without any oxide at all. The results are plotted in the inset of Fig. 18 and the mean ranges given in Table 5. The comparisons show that the ranges are indeed quite sensitive to the oxide. In the channeling case, the ranges are much longer without the oxide. On the other hand, for a 1 nm oxide the results are not too different from the case of 2.2 nm oxide. This shows that even a thin oxide layer is sufficient to scatter many of the ions right at the surface and hence reduce the channeling. In the non-channeling case, the ranges are, on the other hand, slightly shorter without any oxide. This is because the oxide scatters some of the ions into a channel, leading to a longer channeling tail. In both cases, the results for 2.2 nm oxide thickness clearly agree best with the experiments.
Table 5: Comparison of mean ranges $R_{\rm mean}$ for As ranges in Si modelled with the DMol potential for different oxide thicknesses. The error bar is the standard error of the mean. Non-channeling means implantation into directions where channeling is avoided, simulated with the same sample tilt and twist angles as in the experiments.
Case $R_{\rm mean}$ (Å)
No oxide 1 nm oxide 2.2 nm oxide
100 keV non-channeling 712$\pm$2 763$\pm$2 778$\pm$1
15 keV [100] channel 588$\pm$2 376$\pm$ 298$\pm$1
This observation also explains why many of the cases that are not compared to experiments (subsection VI.1) have much stronger channeling effects than the Si cases compared to experiment. In those simulations, the idealized condition of having no surface oxide or angular beam spread maximizes the channeling effects.
VI.2.3 Comparison of results with experiments